OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x - 4. Find f(x) · g(x). x2 - 3x - 28 x2 - 3x - 11 x2 + 3x - 28 x2 + 3x - 11

OpenStudy (anonymous):

Let f(x) = x2 + 3x - 4 and g(x) = x + 5. Find f(x) · g(x). x3 + 3x2 + 16x - 20 x3 + 5x2 + 14x - 20 x3 + 8x2 + 11x - 20 x3 + 9x2 + 19x - 20

OpenStudy (anonymous):

Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x). 2x2 - 4x - 2 2x2 - 4x - 4 2x2 + 6x - 2 2x2 + 6x - 4

OpenStudy (08surya):

x2 + 3x - 28

OpenStudy (anonymous):

Thank you! What about the other two?